### v4

# Displacive Jahn–Teller transition in NaNiO2

Abstract
Below its Jahn–Teller transition temperature, $T_{\mathrm{JT}}$, NaNiO2 has a monoclinic layered structure consisting of alternating layers of edge-sharing NaO6 and Jahn–Teller-distorted NiO6 octahedra. Above $T_{\mathrm{JT}}$ where NaNiO2 is rhombohedral, diffraction measurements show the absence of a cooperative Jahn–Teller distortion, accompanied by an increase in the unit cell volume. Using neutron total scattering, solid-state Nuclear Magnetic Resonance (NMR), and extended X-ray absorption fine structure (EXAFS) experiments as local probes of the structure we find direct evidence for a displacive, as opposed to order-disorder Jahn–Teller transition at $T_{\mathrm{JT}}$. This is supported by ab initio molecular dynamics (AIMD) simulations. To our knowledge this study is the first to show a displacive Jahn–Teller transition in any material using direct observations with local probe techniques.
1 Introduction
In an undistorted octahedrally-coordinated environment, partially-occupied $e_{g}$ orbitals in a transition metal cation would be degenerate and therefore unstable to symmetry-reducing distortions by the Jahn–Teller (JT) effect 1. This lifts the degeneracy of both the $t_{2g}$ and partially-occupied $e_{g}$ orbitals, lowering the energy of the system 1, 2, 3, 4. The resulting octahedral distortion is often a linear combination of two possible van Vleck modes 5, 6: a planar rhombic $Q_{2}$ distortion and a tetragonal $Q_{3}$ elongation/compression [Figure S1]. Experimentally, the distortion is normally dominated by the $Q_{3}$ tetragonal elongation. This JT distortion, and associated orbital ordering, is relevant to many phenomena including unconventional superconductivity in the cuprates 7, 8, magnetic structure through coupling of spin and orbital ordering 9, and ionic mobility 10, 11, and can lead to structural transitions in a material when it is used in a battery electrode 10, 11, 12.
Figure 1: Crystal structures of (a) $C2/m$ monoclinic and (c) $R\bar{3}m$ rhombohedral NaNiO2. (b) and (d) show the Ni-Ni distances (grey for non-elongated; green for elongated) and JT-elongated Ni-O distances (pink) in these two structures. In the rhombohedral structure there are no elongated Ni-Ni or Ni-O bonds.
In a crystalline material, the energy of a JT-distorted system is often minimised when the axes of elongation of neighbouring octahedra are correlated; this is termed a cooperative JT distortion, in contrast to a non-cooperative system in which axes of elongation are randomly distributed 4. Transitions from low-temperature, cooperative JT distortions to a high-temperature state with an undistorted average structure can generally be classified as order-disorder or displacive 13. In the former case, non-cooperative JT distortions persist locally but are averaged out in the bulk structure, whereas in the latter case the local structure is undistorted. This paradigm has been applied to several systems 14, 15, 16, 13, 17. The most well-studied is in the JT-distorted $d^{4}$ Mn3+ perovskite LaMnO3 which shows evidence for an order-disorder JT transition 18, 14, 19, 15; theoretical works 20, 21, 22 and total scattering experiments 16 indicate a transition to a Potts model 20 type of orbital disorder.
The order-disorder/displacive paradigm has also been applied to the layered transition metal oxides with formula $AM$O2 ($A$=alkali metal, $M$=transition metal). 13, 17 This family of materials includes several battery materials such as LiNiO2, NaNiO2, LiMnO2, and NaMnO2, all of which feature transition metal ions with degenerate $d^{7}$/$d^{4}$ $e_{g}$ orbitals which are liable to JT distortions. The $AM$O2 materials have layers of edge-sharing $M$O6 octahedra forming a triangular network of $M$ cations, separated by a layer of octahedrally-coordinated alkali metal ions, in contrast to the perovskites such as LaMnO3 which have corner-sharing octahedra and an approximately cubic cation network. The aristotype of the structure has rhombohedral $R\bar{3}m$ symmetry, but NaNiO2 and each of $\alpha$- and $\beta$-$X$MnO2 ($X$=Li, Na) show cooperative JT ordering 13 resulting in a macroscopic distortion. NaNiO2 exhibits a monoclinic $C2/m$ distortion due to collinear JT-elongated octahedra, which disappears on heating during a first-order transition which onsets at 460 K 23, 24, 25, 26, 27 (see Figure 1) to the aristotype $R\bar{3}m$ structure. LiMnO2 and NaMnO2 exhibit polymorphism and more complex cooperative ordering 13. LiNiO2 is a complicated case, with clear experimental evidence for multiple Ni-O bond lengths 28, 29, 30 but diffraction data are typically modelled with the aristotype structure which does not allow for cooperative JT ordering. There have been a broad array of proposed forms of JT ordering 30, 31 for LiNiO2, with the energetically most-favourable being a zigzag ordering with monoclinic $P2_{1}/c$ symmetry 31, 32, 33, 34, 17. Alternative phenomena for LiNiO2 involving spin- or even charge-disproportionation 32, 34, 35 have also been proposed; these are likely a feature of other nickelates such as AgNiO2 36, 37 or the nickelate perovskites 38, 39, 40. In contrast to a recent theoretical study 13, which concludes that the layered alkali metal transition metal oxides $AB$O2 ($A$=Li, Na; $B$=Ni, Mn) should exhibit an order-disorder JT transition, we have recently found evidence for a displacive transition in LiNiO2 using ab initio Molecular Dynamics (AIMD) and variable-temperature X-ray diffraction (XRD) 17.
[…]
2 Results
2.1 Variable-temperature synchrotron X-ray diffraction.
Figure 2: (a) Representative Rietveld refinement of synchrotron X-ray diffraction data for NaNiO2 at 476.1 K in the mixed-phase regime where the rhombohedral ($R\bar{3}m$) and monoclinic ($C2/m$) phases coexist. (b) Volume per formula unit as a function of temperature, showing the slight increase in unit cell volume of the rhombohedral phase compared with the monoclinic phase. Insets are NaO6 (top left) and NiO6 (top right) octahedral volume with temperature, calculated using VanVleckCalculator; only volumes in the single-phase regions are plotted. Temperature-dependence of the (c) inter-layer distances, $c/3$ for the rhombohedral phase and $c\sin{(\beta)}$ for the monoclinic phase, and (d) intra-layer distances, $a$ for the rhombohedral phase, and $a/\sqrt{3}$ and $b$ for the monoclinic phase. The intra-layer distances correspond to the Ni-Ni and Na-Na distances within the plane. For the monoclinic phase, closed circles are the individual distances, open circles are averaged. In (c), points are experimental values obtained by Rietveld refinement, and solid lines are a linear fit. In (b,c,d) error bars are smaller than data points.
Variable-temperature synchrotron X-ray powder diffraction was performed to observe temperature-dependent changes in the average crystal structure and analysed by Rietveld refinement 41. On heating, the crystal structure of NaNiO2 is monoclinic ($C2/m$) until $\sim$460 K where Bragg peaks associated with the high-temperature rhombohedral ($R\bar{3}m$) phase begin to emerge, indicating this is approximately the onset of the cooperative transition temperature $T_{\mathrm{JT}}$. A mixed-phase regime exists, in which the phase fraction of the monoclinic phase decreases with heating, until the sample becomes entirely rhombohedral around 505 K. These findings are consistent with previous variable-temperature diffraction studies of NaNiO2 24, 25, 26.
Selected results of the Rietveld analysis are shown in Figure 2, with further information in SI Section 2. The volume per formula unit is larger in the rhombohedral than the monoclinic structure [Figure 2(b)]. There is positive thermal expansion in both the monoclinic and rhombohedral phases. In the monoclinic phase, NiO6 octahedral volume is essentially temperature-independent [inset of Figure 2(b)] and the volume increase is entirely driven by NaO6 octahedral expansion. Through the phase transition, there is a discontinuous increase in unit cell volume per formula unit, primarily due to a $\sim$1.2jump in NiO6 octahedral volume (compared with a $\sim$0.6increase for NaO6 octahedra) from the fully monoclinic to fully rhombohedral phases. In the rhombohedral phase both the NiO6 and NaO6 octahedra expand on heating at around the same rate.
Figure 2(c) shows that there is a significant increase in inter-layer distance on heating from the monoclinic to rhombohedral structures. Figure 2(d) shows that the two different intra-layer distances (corresponding to Ni-Ni and Na-Na distances within the plane of the layer) of the monoclinic cell equalise in the rhombohedral cell, due to the increase in cell symmetry, and the distance in the rhombohedral cell is slightly decreased compared with the average distance in the monoclinic cell.
2.2 23Na nuclear magnetic resonance.
A variable-temperature 23Na NMR (VT-NMR) experiment was carried out to investigate changes in the local structure with temperature using both magic angle spinning (MAS) and static measurements. Changes in both of the spectra are observed at $T_{\mathrm{JT}}$ [Figure S37]. At low temperatures a single Na+ environment is observed consistent with the average structure of monoclinic NaNiO2. A second higher chemical shift environment is observed on heating which we ascribe to Na+ in the high-temperature, rhombohedral phase of NaNiO2. There is a limited $T$-range where both peaks co-exist until at higher $T$ only a single peak persists. Both environments have large hyperfine shifts due to the presence of paramagnetic Ni3+ ions, the peak shifting to lower values as the Ni3+ become less paramagnetic at higher temperatures. These results are consistent with the diffraction data and indicate a change in the local Na+ environment at $T_{\mathrm{JT}}$.
Figure 3: (a) Experimental neutron PDFs for NaNiO2 at 293 K, 450 K, and 500 K, in the $r$-range where the nearest-neighbour Ni-O peaks occur. Peaks are labelled for the 293 K dataset in blue text and 500 K dataset in green text. The 500 K data is vertically offset from the low-temperature data for ease of distinguishing the changes compared with the lower-$T$ datasets. Grey dashed horizontal lines occur at $D(r)=0$ for the lower two temperatures (lower) and 500 K (upper). (b,c) Example fits to the PDF and Bragg diffraction data at 500 K using the supercell fitting (experimental data: blue; fitted data: orange; difference pattern; green).
2.3 Neutron pair distribution function.
Total scattering neutron experiments have been performed on NaNiO2. The Bragg data was published previously 27 and is consistent with the synchrotron XRD presented in the previous section.
Pair Distribution Functions (PDFs), presented as $D(r)$ in Figure 3(a), 42 were obtained from the neutron total scattering data at 293 K and 450 K (both monoclinic average structure), and 500 K (rhombohedral average structure). There are two Ni-O peaks at 293 K and 450 K, consistent with the Ni-O bond length splitting due to a Jahn–Teller distortion, and a single Ni-O peak at 500 K, implying that NiO6 octahedra are undistorted by the Jahn–Teller effect at this temperature. This is consistent with the picture from the average structure, as described from the variable-temperature synchrotron data. However, at 500 K the Ni-O peak is highly asymmetric with a tail on the high-$r$ side.
Initial analysis of the PDF data was performed using small-box analysis, also known as real-space Rietveld refinement. Results are presented in Section S3.1 of the SI. The room-temperature data can be fit well with the JT-distorted, monoclinic structure used for Rietveld analysis of the reciprocal space data. The PDF data at 500 K can be fit adequately with both the monoclinic and the rhombohedral (JT-undistorted) structures, with a slightly higher $R_{\mathrm{wp}}$ for the rhombohedral structure (although this is to be expected given the smaller number of free parameters). In the monoclinic cell, the difference between the Ni-O bond lengths is finite but very small at 500 K ($\sim$0.041 Å, compared with $\sim$0.243 Å at 293 K), which is not consistent with the presence of a local Jahn–Teller distortion.
Ultimately the small-box analysis did not fit all features of the PDF data, and in any case small-box analysis with a single unit cell would be insensitive to some types of local Jahn–Teller ordering. To test for the presence of local Jahn–Teller distortions obscured by the peak asymmetry, big-box analysis was performed on the PDF data, in conjunction with the Bragg scattering data.
To this end, a rhombohedral NaNiO2 unit cell with centred orthorhombic setting was prepared, consisting of 6 formula units and obtained via transformation from the normal rhombohedral cell using Equation S5. A $16\times 9\times 3$ supercell of this unit cell was used as the starting point for big box refinements, with lattice dimensions $\sim$45 Å and containing 10368 atoms; see Methods for further details. The atomic positions in this model were then refined against the experimental data. Here, empirical bond valence sum 43 restraints (discussed in Methods) were applied to the atomic positions to ensure physical behaviour, although qualitatively equivalent results are obtained without using these restraints [SI Section S3.9]. The big box refinements were run several times to ensure repeatability [SI Section S3.11].
[…]
Figure 4: Cross sections of the $ab$ plane for a single O-Ni-O layer of the supercell obtained by big box PDF analysis, showing the distribution of elongated ($r>2.1$) Ni-O bonds at (a) 293 K and (b) 500 K. Na ions are hidden for clarity. Ni: grey; O: red.
Figure 5: (a) Ni-O bond distributions at 293 K (top), 450 K (middle), and 500 K (bottom), as obtained via big box analysis of neutron PDF data, for the smallest to largest Ni-O bond in each octahedron; vertical lines are the bond lengths from Rietveld refinement of the neutron data reported in Ref. 27. (b) Histogram of bond length distortion index 44 for each NiO6 octahedron at 293 K, 450 K, and 500 K. (c) Probability function of distortion [Equation 1] at 293 K, 450 K, and 500 K. (b) and (c) are obtained using VanVleckCalculator 6, 45.
Other evidence for the absence of Jahn–Teller distortion in the local structure of rhombohedral NaNiO2 can be obtained by considering two metrics for the octahedral distortion. The bond length distortion index (BLDI) 44, Figure 5(b), quantifies the deviation of bond lengths from the average for a given polyhedron and is often used to quantify Jahn–Teller distortions 46, 47, 27. For a big box model where thermal effects are modelled by deviation of atoms from their average positions (rather than using atomic displacement parameters, as is conventional for Rietveld refinement), we would expect the BLDI to increase with heating in the absence of any electronic or magnetostructural transitions. The average BLDI for all NiO6 octahedra is 0.0547 and 0.0575 (to 3 significant figures) at 293 K and 450 K, respectively, in the monoclinic phase, before decreasing to 0.0458 at 500 K. This is clearly shown by the decreasing distribution of bond lengths in Figure 5(a). In contrast, the average BLDI continuously increases with heating for the NaO6 octahedra, increasing from 0.0260 at 293 K, to 0.0352 at 450 K, and finally 0.0450 at 500 K (see Figure S10). This trend in the BLDI for NiO6 octahedra can be explained by the disappearance of a local Jahn–Teller distortion at $T_{\mathrm{JT}}$. We can also quantify the distortion using the parameter $\rho_{0}$, where $\rho_{0}^{2}=Q_{2}^{2}+Q_{3}^{2}$. Figure 5(c) shows $P(\rho_{0})$, the probability of octahedra having a given value of $\rho_{0}$, defined as:
$P(\rho_{0})=\frac{f(\rho_{0})}{2\pi\rho_{0}}$ (1)
where $f(\rho_{0})$ is a histogram of $\rho_{0}$ for all NiO6 octahedra. This parameterisation shows maximum probability at $\rho_{0}\approx 0$ at 500 K, compared with large non-zero magnitudes at 293 K and 450 K, which also indicates the absence of a local Jahn–Teller distortion in the rhombohedral phase. This result that the most probable $\rho_{0}$ substantially decreases at 500 K, compared with lower temperatures, is resilient to other ways of running the big box refinement, for instance with different starting configurations [SI Section 3.7], changes in the BVS restraints [SI Section S3.9], and multiple runs with the same starting parameters [SI Section 3.11].
2.4 X-ray absorption spectroscopy.
Figure 6: (a) XANES at the Ni K edge at various temperatures, where each spectrum is offset by an arbitrary amount for clarity; inset shows $d\mu/dE$ for the 298.15 K and 533.15 K, with an asterisk marking the feature on the rising edge. (b) EXAFS data at the Ni K edge, with heating, for NaNiO2 at various temperatures, with the solid line representing the fit with a {2,2} model consisting of two Ni-O bond lengths and two Ni-Ni interatomic distances. The feature around 3.4 Å is the Ni-Na atomic distance, which was not included in any of the fitting. (c) Short and long Ni-O bond lengths obtained from fitting the {2,2} and {2,1} models, to the EXAFS data; Ni-O lengths from fitting the {1,1} model are also shown. (d) Figures of merit for fits performed with the {2,0} and {1,0} models, reduced $\chi^{2}$ and Bayesian Information Criteria (BIC), in the vicinity of the Ni-O bond only.
XAS at the Ni K edge was performed, with analysis focusing on the extended X-ray absorption fine structure (EXAFS) in real-space. This data, $\chi(r)$, is analogous to a partial PDF of Ni-O bond lengths, but with the significant difference that there is a phase shift resulting in peaks in $\chi(r)$ being down-shifted by around 0.5 Å. It is therefore a useful supplement to the PDF data. The room-temperature $\chi(r)$ data presented here resembles the EXAFS on NaNiO2 reported previously 48.
Both the X-ray absorption near edge structure (XANES) and EXAFS change on heating through $T_{\mathrm{JT}}$ [Figure 6]. In the XANES, although there is no change in peak position (indicating that the Ni oxidation state does not change from Ni3+) there is a pre-edge feature around 8335 eV in $d\mu/dE$ which is more prominent for the $T<T_{\mathrm{JT}}$ regime than the $T>T_{\mathrm{JT}}$ regime. This pre-edge is likely due to quadrupole forbidden transitions, and will be less prominent in higher symmetry Ni environments. This observation is therefore consistent with the absence of a local Jahn–Teller distortion at $T>T_{\mathrm{JT}}$.
The EXAFS data, Figure 6(b), also shows a significant change on heating. The two peaks corresponding to two Ni-Ni interatomic distances at $\sim$2.5 Å when $T<T_{\mathrm{JT}}$ merge into a single peak in the $T>T_{\mathrm{JT}}$ regime, consistent with the single Ni-Ni interatomic distance in the rhombohedral structure. The peak at $\sim$1.5 Å corresponds to the Ni-O bond, which exhibits a far more subtle change in shape through the Jahn–Teller transition than the Ni-Ni peak, and fitting to the data is required to examine the changes in local structure.
Models were obtained using FEFF 49 and refined against the $\chi(r)$ data over different ranges. Two different R(Å) ranges were considered; one with just the Ni-O shell R(Å) = 0.5 – 2.0 Å, and the other R(Å) = 0.5 – 3.1 Å to also include the Ni-Ni interatomic distances. Models with both 1 or 2 Ni-O and Ni-Ni distances were considered and we describe these using curly braces enclosing the comma-separated values for the number of lengths for each bond; for instance a model with a single Ni-O distance in the reduced range is described as {1,0}, whereas a model with two Ni-O bond lengths and two Ni-Ni bond lengths is described as {2,2}. Five models were used, over both ranges: {1,0}, {2,0}, {1,1}, {2,1}, and {2,2}. Models in which the first number is 1 correspond to JT-undistorted, and models where the first element is 2 are compatible with a JT distortion.
[…]
Over the larger range (0.5 Å to 3.1 Å) fitting was performed using models {2,2}, {2,1}, and {1,1}. When $T<T_{\mathrm{JT}}$, the {1,1} and {2,1} models perform poorly as they cannot reproduce the Ni-Ni splitting observed experimentally, Figure 6(b). The temperature-dependence of the figures of merit, Figure S30, indicate that below $T_{\mathrm{JT}}$, {2,2} is the favoured model, consistent with diffraction. Above $T_{\mathrm{JT}}$, all three models qualitatively fit the data well, but the two figures of merit support different models; BIC favours the undistorted model {1,1} whereas $r\chi^{2}$ the {2,2} model.
[…]
2.5 Ab initio molecular dynamics.
Figure 7: (a,c,d,f) Snapshots of a NiO6 layer as obtained from AIMD simulations with $C2/m$ (a,c) and $R\bar{3}m$ (d,f) starting configurations, at various temperatures. Ni: grey; O: red. Ni-O bonds above 2.1 Å are shown. (b,e) Van Vleck $E_{g}(Q_{2},Q_{3})$ diagrams as a function of temperature for both configurations, showing the displacive transition; note that a rotation of 120 from the $Q_{2}=0$ line means a change in the axis of elongation. Analogous figures based on the PDF big box analysis can be seen in Figure S11.
Figure 8: (a) Calculated Ni-O pair distribution functions from AIMD; these are convolved with experimental $Q_{\mathrm{max}}$ in Figure S36. (b) The probability function of $\rho_{0}$ [Equation 1] from AIMD, for the cells with the $C2/m$ starting structure. (c) Ni-O bond lengths at two temperatures for AIMD runs from the $C2/m$ starting structure.
Ab initio molecular dynamics simulations of NaNiO2 simulation cells starting with a collinear ordering of Jahn–Teller distortions, in a $3\times 3\times 3$ supercell (216 ions) expanded from the monoclinic cell, were performed at temperatures between 34 and 1207 K [Figure 7(a,c)]. Both the cell shape and volume were allowed to change. At low temperatures, the AIMD trajectories show thermal vibrations but the JT distortions persist, and remain collinear. On heating, the distortions decrease in magnitude and are no longer aligned to a single axis. At high temperatures, the octahedra approach a JT-undistorted state.
The temperature-dependence of the Jahn–Teller distortions is quantified by looking at the Ni-O PDF [Figure 8(a)] and the probability function of $\rho_{0}$ [Figure 8(b)]. At low temperatures (34 K), the PDF shows distinct peaks corresponding to short and long Ni-O bonds with a large, positive $\rho_{0}$. On heating (972 K, 1207 K) the two peaks in the PDF merge to a single broad asymmetric peak. $\rho_{0}$ significantly decreases, with the highest probability at $\rho_{0}=0$. This can also be seen in the heat map distributions in $E_{g}(Q_{2},Q_{3})$ shown in Figure 7.
However, at these high temperatures there remains a residual distortion evident from a small asymmetry in the Ni-O PDF distribution, with a preference for elongation along the collinear axis [Figure 7(b) 1207 K]. We attribute this to the finite simulation cell size enforcing a periodicity in the structure that may cause longer-range correlations of the NiO6 octahedra. To explore this, simulations were run from a similarly-sized starting configuration ($4\times 4\times 1$ expansion of the rhombohedral unit cell; 192 ions) without JT distortions, and hence without inter-octahedral correlations to begin with, cooling the sample from 1000 K. The same underlying behaviour is observed as seen in the simulations starting from JT distorted simulation cells, i.e. distortions are observed at $T<T_{\mathrm{JT}}$ that vanish at $T>T_{\mathrm{JT}}$ and (re)emerge at $T<T_{\mathrm{JT}}$. At ca. 1000 K, the NiO6 octahedra are fully isotropic [Figure 7(e) 1000 K]. On cooling, Jahn–Teller distortions emerge. These are row-ordered [Figure 7(e) 300 K], i.e. arranged collinearly within a row, rather than across the simulation cell, as predicted previously 33. The rows show a preference for two out of the three possible directions of the distortions (i.e. $\arctan{(Q_{2}/Q_{3})}=0$ or $32\pi/$ 3 but never $34\pi/$ 3), resulting in two clusters of data points in the van Vleck plot at low temperatures [Figure 7(e) 50 K]. This is likely a consequence of the very small domain sizes, resulting presumably from the rapid cooling on the timescale of the AIMD simulations (a few picoseconds), not allowing the system to arrange into a macroscopically collinear ground state. We note that in the transition regime for both simulation cells, the octahedral distortions are elongated along two out of the three octahedral axes [Figure 7(b) 972 K,(e) 300 K], suggesting the transitions between the collinearly distorted phase with elongations along one axis and the displacive phase with random fluctuations along all axes may occur through an intermediate state with distortions occurring preferentially along two axes. If cooled slowly, the distortions are expected to order collinearly, whereas the rapid quenching of the AIMD simulation freezes the distortions in.
The calculated transition temperature between the JT-distorted state and the JT-undistorted state depends on the initial configuration of the simulation cell, and this does not change when the length of the AIMD runs is extended further. On heating the distorted simulation cell, the onset of the JT-transition $T_{\mathrm{onset}}$ is found to occur at around 586 K, i.e. at a slightly higher temperature than the experimental onset temperature obtained from synchrotron X-ray diffraction, 460 K. On cooling the undistorted simulation cell, the onset of the transition $T_{\mathrm{onset}}$ is below 300 K, i.e. slightly lower than the experimental onset temperature. A closer look at these differences in $T_{\mathrm{onset}}$ reveals that there are two factors causing the deviations. First, as the transition is a first-order transition, hysteric behaviour is expected, i.e. the transition temperature is expected to differ between heating and cooling, and, albeit to a smaller extent, hysteresis has been observed experimentally, too. 25, 26 Second, the AIMD simulations starting from the distorted and undistorted simulation cells result in different domain sizes, affecting the transition temperatures further. The cooperative collinear distortions of the distorted simulation cell correspond to a scenario of infinitely large domains which shift the transition temperature to its maximum value. When cooling the undistorted simulation cell, very small domains form [Figure 7(e) 300 K] resulting in short-range correlations between octahedral distortions. The transition temperature obtained from cooling the undistorted simulation cell therefore constitutes a lower cutoff to the transition temperature from the JT-undistorted phase to the distorted phase. The AIMD simulations thus predict the transition from the JT-distorted to the JT-undistorted phase to occur at $T<586$ K, and the transition from the JT-undistorted to the JT-distorted phase at temperatures $T>300$ K. The onset temperature derived for the transition to the JT-undistorted phase based on synchrotron XRD of samples of intermediate domain sizes in this study, 460 K, lies approximately in the middle of the predicted temperature window.
3 Discussion
Our studies of the local Ni3+ environment, using neutron PDF and Ni K edge EXAFS, suggest the absence of local static Jahn–Teller-distortions for $T>T_{\mathrm{JT}}$. Further, while a dynamic Jahn–Teller distortion 4 has previously been suggested for LiNiO2 51, 52, 53, by analogy with similar proposals for LaMnO3 54, such a dynamic Jahn–Teller distortion would also be resolvable via our neutron PDF and EXAFS measurements. These observations suggest that the JT transition in NaNiO2 is better characterised as displacive rather than order-disorder 13. This is different to the observations of an order-disorder transition in LaMnO3 18, 15, 14, 19, and the proposed order-disorder transition for NaNiO2 13. To our knowledge, only in LiNiO2 17 has prior evidence for a displacive Jahn–Teller transition been put forward, using AIMD and inferences from changes in strain.
As in previous reports we observe a first order phase transition in NaNiO2 with the monoclinic and rhombohedral phases coexisting over a finite temperature range. The observed discontinuous increase in unit cell volume from the monoclinic to the rhombohedral structures [Figure 2] is also consistent with a first-order phase transition. The positive thermal expansion through the transition is consistent, via the Clausius-Clapeyron equation, with our previous finding 27 of $dT_{\mathrm{JT}}/dP>0$. We note that in the displacive transition in SrTiO3 55 a positive $dT_{\mathrm{a}}/dP$ is also observed. In contrast, a negative thermal expansion is found at the transition temperature in LaMnO3 56 and ferroelectric PbTiO3 57, 58, both order-disorder transitions, with $dT_{c}/dP<0$ 59, 60. The displacive transition may therefore be related to the sign of $dT_{\mathrm{JT}}/dP$.
In NaNiO2, the high-temperature phase will be enthalpically unfavourable and hence the phase transition will be entropically-driven, according to the free energy $\Delta G=\Delta H-T\Delta S$, where $\Delta H$ is the enthalpy change, $T$ is temperature, and $\Delta S$ is the entropy change. There are two relevant contributions to the entropy: configurational and vibrational. We find the configurational entropy of orbital disorder for LiNiO2 and NaNiO2 to be subextensive (see Appendix A), meaning it does not increase with increasing system size 61. A subextensive configurational entropy has also been found for an Ising model on an elastic triangular lattice 62. Consequently the entropy term is dominated by dynamic vibrational effects which disfavour local Jahn–Teller distortions. This is likely a contributing factor in the displacive JT transition we have observed, and is in contrast to high-temperature LaMnO3 where, by analytic and Monte-Carlo calculations using the Potts model, there is extensive configurational entropy of orbital disorder 20, 21 and hence an order-disorder transition.
An alternative hypothesis we considered for the disappearance of Jahn–Teller distortions in high-temperature NaNiO2 was the possibility for delocalisation of the electrons resulting in a change in the energy landscape of the system. This is by analogy to the insulator-to-metal transition (IMT) previously observed in the perovskite nickelates 63, 64. However, while variable-temperature conductivity, $\sigma(T)$, data from Ref. 65 [Figure S6], shows an increase in conductivity by several orders of magnitude on heating, there is no change of sign of $d\sigma(T)/dT$. This is similar to the conductivity data for LaMnO3 54, which would suggest the electronic conductivity is not closely related to the order-disorder/displacive nature of the transition.
There are a number of other differences between NaNiO2 and LaMnO3 which could affect the Jahn–Teller transition. For instance, the edge-sharing octahedral connectivity in NaNiO2 results in much larger intra-octahedral angular distortion compared with the corner-sharing octahedra in LaMnO3. This difference in connectivity is a consequence of the triangular Ni network in NaNiO2. Given the constraint that axes of elongation of adjacent JT-distorted Ni atoms cannot point at the same O anion, there are no possible long-range JT-disordered configurations 30, which means that this triangular Ni network cannot host a Potts-type Jahn–Teller disorder such as that proposed for LaMnO3 20, 21, 16. Another difference is that $d^{4}$ Mn3+ is generally found to be JT-distorted, whereas nominally $d^{7}$ Ni3+ is susceptible to other types of electronic configuration such as charge disproportionation 38, 39, 36, 37, 40; this may be due to the smaller electronegativity difference between Ni-O, compared with Mn-O, which has been argued to cause charge-transfer insulating 66 behaviour in many nickelates 67, 64, 68. Attempting to deconvolute the role of these different factors will require further study on the transition in other Jahn–Teller-distorted materials, along with theoretical and computational work.
Whether or not a JT transition is displacive or order-disorder may be important in some battery cathode materials. This will be particularly apparent if the JT transition occurs below room-temperature, as it does in, for instance, the NaNixCo1-xO2 solid solution for $x\leq 0.8$ 65 and the spinel LiMn2O4 69. In these cases, the presence or absence of local JT distortions will result in different ionic mobility, and will impact the crystallographic changes that occur with electrochemical cycling. The implications of this on electrochemistry should be studied in future works.
4 Conclusion
We have shown that the Jahn–Teller transition in NaNiO2 is displacive in character, with the absence of local JT-distortions or orbital ordering in the NiO6 octahedra at $T>T_{\mathrm{JT}}$ ($T_{\mathrm{JT}}$ being the cooperative Jahn–Teller transition temperature). To our knowledge this is the first time that local probes of structure have shown a displacive Jahn–Teller transition, though order-disorder transitions have been reported for other JT-distorted perovskite LaMnO3 18, 19, 15, 14.
Our findings are not consistent with previous computational studies on NaNiO2, which predicted an order-disorder transition with persisting local JT distortions 13, but our own ab initio molecular dynamics calculations support our experimental findings. This complementarity between local probe experiments and simulation provides further support for our earlier ab initio molecular dynamics study on LiNiO2 which also predicted a displacive JT transition 17.
We found that the configurational entropy of orbital disorder is subextensive in layered transition metal oxides such as NaNiO2, and contrast this to the extensive configurational entropy 20 in the case of perovskites such as LaMnO3. Further theoretical work is required to better understand these different behaviours.
Methods
4.1 Sample synthesis.
Samples were prepared by solid state synthesis. Na2O2 (Alfa Aesar; 95%) and NiO (Alfa Aesar; 99.995%) were mixed and pelletised in a 1.05:1 molar ratio of Na:Ni, with excess Na to account for Na-loss during heating. The pelletised precursor mixture was placed in an alumina crucible and heated to 973 K for 70 hrs in a tube furnace under a constant flow of O2. O2 was maintained throughout the heating and cooling processes. The sample showed a colour change from light green (the NiO/Na2O2 precursor mixture) to black, indicating successful synthesis. To prevent reaction with moisture, the sample was stored and handled in an inert Ar-atmosphere. The same sample was previously studied in Ref. 27, where electron microscopy and Rietveld 41 analysis of laboratory X-ray diffraction and neutron diffraction data are presented.
4.2 Synchrotron X-ray diffraction.
Variable-temperature, ambient-pressure synchrotron X-ray diffraction was performed using the I11 instrument 70, 71 at Diamond Light Source ($\lambda=0.824110$) using the Mythen-2 position-sensitive detector, with a data collection time of $\sim$10 seconds. The sample was contained in a 0.5 mm diameter glass capillary, sealed inside a glovebox with epoxy (Loctite Double Bubble). The sample was heated at a rate of 12 K/min from 322 K to 796 K, with periodic measurements.
Data were analysed by sequential Rietveld refinement 41 using the software package Topas 7 72. A pseudo-Voigt peak function was used, and background was fitted using a Chebyshev polynomial with 20 terms. A small correction for preferred orientation was applied, using the March-Dollase model 73, 74.
4.3 Nuclear magnetic resonance.
Samples for variable-temperature nuclear magnetic resonance (VT-NMR) were packed into 4 mm ZrO2 magic angle spinning (MAS) rotors in an Ar filled glovebox, and fitted with ZrO2 caps. 23Na chemical shifts were calibrated using solid NaCl as an external secondary reference (7.21 ppm relative to 1 M NaCl(aq) at 0.0 ppm). 23Na MAS NMR spectra were acquired using a Bruker Avance IIIHD spectrometer (v0[1H] = 500.13 MHz, v0[23Na] = 132.46 MHz, v0 [207Pb] = 104.26 MHz) spectrometer, with a Bruker 4 mm HX probe and 14 kHz MAS. Projection magic-angle turning and a phase-adjusted sideband separation (pj-MATPASS) pulse sequence was used to obtain the isotropic 23Na spectra 75.
Static VT-NMR samples were measured in the same spectrometer using a Bruker HX static probe. A Hahn-echo pulse sequence with $\pi/4$ = 2.05 $\mu$s optimized on solid NaNiO2 was used. Temperature was calibrated through a NMR shift thermometer compound Pb(NO3)2, based on the known temperature-dependence of the isotropic chemical shift of Pb(NO3)2 76. 207Pb NMR spectra were acquired with the same probe at the same temperature values, using a Hahn-echo pulse sequence. The isotropic chemical shift values for Pb(NO3)2 were obtained by fitting the data using the SOLA fitting program within Topspin 4.1.4.
4.4 Neutron total scattering.
Variable-temperature total-scattering neutron diffraction was performed on the NOMAD instrument 77 at the Spallation Neutron Source, Oak Ridge National Laboratory, USA. NaNiO2 was stored under Ar in a sealed NMR tube (2 cm sample height, 5 mm outer diameter) for the measurements. Heating was performed using a furnace, at a rate of 5K per minute. The sample was measured on heating at 293 K, 450 K, 500 K, and after cooling at 316 K. The neutron diffraction data and the refined crystal structure was previously published in Ref. 27. In the Fourier transform to real space, a sliding $Q_{\mathrm{max}}$ as a function of $r$ is used, with $Q_{\mathrm{max}}=40$-1 at low-$r$ and gradually decreasing $Q_{\mathrm{max}}$ beyond that, with $Q_{\mathrm{max}}=20$-1 above $r=25$.
4.5 Analysis of pair distribution function.
Small-box pair distribution function analysis, also known as real-space Rietveld refinement, was performed using Topas 6 72. Data were fitted in the range 1.0 Å to 10 Å in real space only. The starting structures were the normal $C2/m$ and $R\bar{3}m$ structures used commonly in the literature for the low- and high-temperature phases of NaNiO2.
Pair distribution function analysis was performed via a “big box” approach used Topas 7 72, following a modified version of the method introduced in Ref. 78. A supercell model was refined against both the Bragg data from NOMAD bank 4 (as defined in Ref. 77) and the experimental neutron pair distribution function. In this refinement, atomic coordinates were refined until convergence following the Rietveld algorithm 41. During initial iterations, restraints were applied to prevent large atomic movements, but these were removed in later stages of the refinement. Upon convergence, each atom was randomly shifted by a distance less than or equal to 0.1% of the unit cell size ($\sim$0.045 Å) in each spatial dimension to help minimisation. This typically occurred several hundred times. The final configuration was the converged structure with the best fit quality. Our approach differs from the more commonly-used approach in RMCProfile 79. Data were fitted in the range 0.5 Å to 20 Å in real space.
A $16\times 9\times 3$ supercell of a pseudo-orthorhombic cell (space group: $P1$; $a\approx 2.846$; $b\approx 5.321$; $c\approx 15.703$; $\beta=\gamma=90^{\circ}$; $\alpha=90^{\circ}$ for the 500 K supercell and $\sim$88 for the $T<T_{\mathrm{JT}}$ supercells) was used for the refinements. This NaNiO2 unit cell was obtained via transformation from the rhombohedral cell using Equation S5 in Supplementary Information. The supercell has edge lengths $\sim$45 Å and contains 10368 atoms. In this analysis, the thermal parameters for all atoms were fixed at low values ($B_{\mathrm{iso}}=0.01$2) such that thermal effects are modelled by the deviation of each site from its ideal position. To ensure that chemically reasonable models were obtained, soft restraints were included. The relative weighting of restraints against data was set to ensure refinements were not dominated by restraints. The SI includes refinements performed in the absence of restraints to check their influence on the conclusions (Section S3.9). The main soft restraint was the calculated bond valence sum 43 (as an analog for oxidation state), $V_{\mathrm{calc}}$, of cations deviates from its expected value (Na: +1; Ni: +3), an established approach 80 in PDF analysis 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 78, 100, 101, 102, 103. Here, $V_{\mathrm{calc}}$ is calculated as follows:
$V_{\mathrm{calc}}=\sum_{i=1}^{6}\exp{\left[\frac{r_{0}-r_{i}}{B}\right]}$ (2)
where the empirical parameters $r_{0}$ and $B$ are based on the cation and anion species, and $r_{i}$ are the bond lengths between the metal atom and the $i$th oxygen. In this work, $B=0.37$ and $r_{0}^{\mathrm{Na}}=1.672$ and $r_{0}^{\mathrm{Ni}}=1.7335$; these values were selected to give average bond lengths of 2.335 Å and 1.99 Å for NaO6 and NiO6 octahedra respectively, to match the room-temperature average structure.
Since the $P1$ space group has a floating origin, a restraint was applied to keep the average shift of Ni sites close to zero.
The supercells obtained from refinement were then analysed using VanVleckCalculator 6, 45, a Python 3 104 code which is based on PyMatGen 105.
4.6 Ni K edge X-ray absorption spectroscopy.
X-ray absorption spectroscopy (XAS) measurements were performed on the BM23 instrument 106 of the European Synchrotron Radiation Facility (ESRF), France. Powdered NaNiO2 was mixed homogeneously with a powdered boron nitride binder ($\sim$16 mg:$\sim$70 mg NaNiO2:BN) and pressed into a 13 mm pellet in a dry, oxygen-free nitrogen environment. The pellet was cut and placed in a sealed sample holder under flowing helium. Temperature control was achieved using a resistive heater, and at each temperature the sample was left for 6 minutes to thermally equilibriate. At each temperature the sample was measured ten times to ensure reproducibility, and an average spectrum produced. X-ray absorption was measured in the vicinity of the Ni K edge ($\sim 8.3$ keV).
X-ray absorption was calculated using the Beer-Lambert law 107 comparing X-ray intensity in ionisation chambers before and after transmission through the sample, and a third ionisation chamber measured the absorption through Ni foil as a reference. Data normalisation was performed using the Python-based 104 Larch 108 package. For the normalisation of the data, a pre-edge range between -350 eV and -45 eV from the edge centre was fit with a straight line, and the post-edge region between 200 eV and 1250 eV was fit with a quadratic polynomial. Data were transformed from $E$-space to $k$-space via the transform $k=\sqrt{\frac{2m_{e}}{\hbar^{2}}\left(\hbar\omega-E_{0}\right)}$, where $m_{e}$ is the mass of the electron, $E_{0}$ is the edge energy, and $\omega$ is the frequency of the measured photon. Finally, the EXAFS function $\chi(k)$ was obtained by fitting and subtracting a spline to the EXAFS data in $k$-space, where $R_{\mathrm{bkg}}$ was set to 1 Å and the spline was fit within a $k$-range of 0 Å-1 to 18.25 Å-1, with a $k^{2}$-weighting. For the Fourier transform of EXAFS from $k$-space to $R$-space, a $k$-range of 2.5 Å-1 to 15 Å-1 was used, and to minimise Fourier ripples from the cutoffs in $k$-space, a Kaiser-Bessel window with end width $dk=7$-1 was applied to the data in $k$-space before the transformation.
Calculated EXAFS data were obtained by multiple-scattering path expansion using FEFF 49, as implemented in Larch in Python. In this model, data is fit using the EXAFS equation over each set of equivalent paths $j$, defined as follows:
$\chi(k)=S_{0}^{2}\sum_{j}\frac{N_{j}|f_{j}(k)|}{kR_{j}^{2}}\sin{\left[2kR_{j}+\phi_{j}(k)\right]}\\ \times\exp{\left[-\frac{2R_{j}}{\lambda_{j}(k)}\right]}\exp(-\sigma_{j}^{2}k^{2})$ (3)
where FEFF is used to calculate scattering amplitude $|f_{j}(k)|$, photoelectron mean free path $\lambda_{j}(k)$, and phase $\phi_{j}(k)$ for all paths. $N_{j}$ is the number of equivalent paths for each $j$, which is fixed. The amplitude reduction factor $S_{0}^{2}$ was kept fixed at a value obtained by fitting against the Ni foil reference, $S_{0}^{2}=0.79286024$. The path distance $R_{j}$ and Debye-Waller factors $\sigma_{j}$ for each path were refined during model fitting. Edge position $E_{0}$ was also freely refined.
4.7 Ab initio molecular dynamics.
AIMD simulations were performed according to the Generalized Gradient Approximation (GGA) 109 and the projector augmented wave method (PAW) 110, as implemented in the Vienna Ab Initio Simulation Package (VASP) 111, 112. The plane-wave energy cutoff was set to 500 eV. Supercells are expansions of the monoclinic ($3\times 3\times 3$; 216 ions) and rhombohedral ($4\times 4\times 1$; 192 ions) unit cells, respectively. Simulations were performed at the $\Gamma$ point and checked against calculations with a $2\times 2\times 2$ Monkhorst-Pack 113 k-point grid for consistency. The convergence criteria for the electronic and ionic relaxations were set to $10^{-6}$ eV and $5\cdot 10^{-3}$ eV/Å, respectively.
For Ni, the $4s^{2}3d^{8}$ electrons were treated as valence electrons. To account for the strongly correlated d electrons, a rotationally invariant Hubbard U parameter 114 of $U_{\mathrm{eff}}=6$ eV was selected, which was used successfully in previous studies of layered oxide cathodes, both for 0 K DFT calculations and finite temperature AIMD simulations 115, 116, 17. For oxygen, the $2s^{2}2p^{4}$ electrons were considered as valence electrons.
AIMD simulations were performed for the isothermal-isobaric ensemble (NpT, constant pressure, particle number, and temperature) at zero pressure, allowing the cell shape and volume to change. A Langevin thermostat was used with friction coefficients set to zero to minimise impact on the lattice vibrations. Trajectories were run for 3 ps at timesteps of 1 fs. Following an equilibration period of 1 ps, snapshots were sampled every 2-5 fs, using Ovito 117, for the analysis of the distortions.
Appendix A: proof of subextensive configurational entropy for orbital disorder
Here, a mathematical proof is presented that, given a set of assumptions, there is subextensive configurational entropy of orbital disorder in layered oxide materials based on a triangular array of JT-active cations. We define subextensive as meaning that the configurational entropy does not increase with increasing system size.
This relies on three basic axioms:
  1. 1.
    The effect on orbital ordering of inter-layer interactions is negligible.
  2. 2.
    Each oxygen anion participates in precisely one Jahn–Teller-elongated bond.
  3. 3.
    Jahn–Teller distortions are purely of the $Q_{3}$ tetragonal elongation type.
[…]
Configurational entropy per layer is given by:
$S_{\mathrm{config}}^{\mathrm{layer}}=\ln{\left[W_{\mathrm{layer}}\right]}$ (A.1)
where $W_{\mathrm{layer}}$ is the number of configurations within the layer. Note we define entropy in units of Boltzmann’s constant $k_{\mathrm{B}}$.
Given Axiom 2, there are only two possible alignments of the elongated axes between neighbouring octahedra. Either the axes of elongation are parallel (“A”) or angled 120 apart (“B”). There are three rows of Ni sites within a layer, separated by an angle 120. In a monoclinic cell such as in NaNiO2 at $T<T_{\mathrm{JT}}$, all interactions would be of the “A” type. If there were true orbital disorder, each chain would be a random combination of these two possible configurations “A” and “B”, for instance ”AABBABAAABAABBABABB”. If L is the length of such a chain (divided by Ni-Ni distance), then we have the following number of configurations:
$W_{\mathrm{layer}}=3\times 2^{L}=3\times 2^{A^{1/2}}$ (A.2)
where the $3$ factor occurs because there are three directions where chains occur, and we relate $L$ to an arbitrary area $A$ via $L=A^{1/2}$.
Given that $N$ is proportional to $A$, where $N$ is the number of Jahn–Teller-elongated axes within an area of the single layer, we can use $N=\alpha A$ and substitute this into the equation for configurational entropy as follows:
$S_{\mathrm{config}}^{\mathrm{layer}}=\ln{\left[3\times 2^{\alpha^{\prime}N^{1/2}}\right]}=C+N^{1/2}\ln{\left[2\right]}$ (A.3)
where we have used $\alpha^{\prime}=\alpha^{1/2}$ and used constant $C=\ln{\left[3\right]}+\alpha^{\prime}\ln{\left[2\right]}$.
If we then calculate configurational entropy per NiO6 octahedron for a large system, as $N\rightarrow\infty$, $C/N\rightarrow 0$ and so:
$\frac{S_{\mathrm{config}}^{\mathrm{layer}}}{N}\approx N^{-1/2}\ln{\left[2\right]}$ (A.4)
Therefore, as $N\rightarrow\infty$, so does $\frac{S_{\mathrm{config}}^{\mathrm{layer}}}{N}\rightarrow 0$. This indicates that the configurational entropy of orbital disorder is subextensive.
Data availability
Data from the BM23 instrument at the European Synchrotron Radiation Facility is available at doi:10.15151/ESRF-ES-962076745 118. All other data is available in the University of Cambridge repository at doi.org/10.17863/CAM.112349 119.
Supplementary Information
A document containing Supplementary Information to this article is available. The Supplementary Information contains: A description of the Van Vleck modes. Further analysis of supercell outputs of big box PDF analysis, and results from various different starting configurations. Expanded AIMD results. NMR data. Tabulated results of all analysis techniques.
Acknowledgements
The authors acknowledge Oak Ridge National Laboratory, a United States Department of Energy Office of Science User Facility, for use of the NOMAD instrument at the Spallation Neutron Source (experiment IPTS25164). We acknowledge the European Synchrotron Radiation Facility for provision of beam time on BM23 (experiment CH6437). We acknowledge I11 beamline at the Diamond Light Source, UK, for the synchrotron XRD measurement done under BAG proposal (the data presented in this work under CY34243; essential preliminary data from CY28349). Calculations were performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick (EP/T022108/1).
We would like to thank Dr Euan N. Bassey, Lucy Haddad, Dr Anastasia Yu. Molokova, Dr Chloe C. Coates, Dr Farheen N. Sayed, Dr Gheorghe-Lucian Pășcuț, and Dr João Elias F. S. Rodrigues for useful discussions.
Crystal structure figures were prepared using Vesta-3 120 and graphs were prepared using MatPlotLib 121.
Funding
This work was supported by the Faraday Institution (FIRG001, FIRG017, FIRG024, FIRG060). L.A.V.N-C acknowledges a scholarship EP/R513180/1 to pursue doctoral research from the UK Engineering and Physical Sciences Research Council (EPSRC) and additional funding from the Cambridge Philosophical Society. J.M.A.S. acknowledges support from the EPSRC Cambridge NanoCDT, EP/L015978/1. A.L.G. acknowledges European Research Council (ERC) funding under grant 788144.